# Convergence and long-time behavior of finite volumes for a generalized Poisson-Nernst-Planck system with cross-diffusion and size exclusion

### v1

@@ 3.2. Energy dissipation at the discrete level
Next proposition is about the thermodynamical consistency of our scheme and the decay of a discrete counterpart of the free energy.
Proposition 3.5.
@@
(41) $\mathcal{H}_{\mathcal{T}}^{n}=\sum_{K\in\mathcal{T}}m_{K}H(U_{K}^{n})+\frac{\lambda^{2}}{2}\sum_{\sigma\in\mathcal{E}}a_{\sigma}(\phi_{K}^{n}-\phi_{K\sigma}^{n})^{2}+\lambda^{2}\sum_{\sigma\in\mathcal{E}^{D}}a_{\sigma}\phi_{\sigma}^{D}(\phi_{K}^{n}-\phi_{\sigma}^{D}),$
@@
$\mathcal{D}_{\mathcal{T}}^{n}=\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}F_{i,K\sigma}^{n}(\mu_{i,K}^{n}-\mu_{i,L}^{n})$
@@
(43) $\mathcal{D}_{i,\sigma}^{n}:=F_{i,K\sigma}^{n}(\mu_{i,K}^{n}-\mu_{i,L}^{n})\geq 0,\qquad\forall i=1,\dots,I,\;\sigma=K|L\in\mathcal{E}_{\text{int}},$
whence the nonnegativity of $\mathcal{D}^{n}$. As $u_{0,K}^{n}>0$ for all $K\in\mathcal{T}$ and $n\geq 1$, and as $y\mapsto e^{y}$ is strictly increasing, one gets that $\mathcal{D}_{i,\sigma}^{n}=0$ iff $\mu_{i,K}^{n}=\mu_{i,L}^{n}$.
Define by $\mu_{i,K}^{n}=\log\left(\frac{u_{i,K}^{n}}{u_{0,K}^{n}}\right)+z_{i}\phi_{K}^{n}=\log(w_{i,K}^{n})$ the electrochemical potential of species $i$, then multiplying the discrete conservation law (25) by $\tau^{n}\mu_{i,K}^{n}$, and summing over $i=1,\dots,I$ and $K\in\mathcal{T}$ provides thanks to discrete integration by parts
(44) $\mathcal{A}_{\mathcal{T}}^{n}+\mathcal{B}_{\mathcal{T}}^{n}+\tau^{n}\mathcal{D}_{\mathcal{T}}^{n}=0,$
@@
${=}$ $\sum_{i=0}^{I}\sum_{K\in\mathcal{T}}\left(u_{i,K}^{n}-u_{i,K}^{n-1}\right)\log\left({u_{i,K}^{n}}\right)m_{K},$
and
@@
${=}$ $\lambda^{2}\sum_{K\in\mathcal{T}}\phi_{K}^{n}\sum_{\sigma\in\mathcal{E}_{K}}a_{\sigma}\left(\phi_{K}^{n}-\phi_{K}^{n-1}-(\phi_{K\sigma}^{n}-\phi_{K\sigma}^{n-1})\right).$
Then we deduce from the convexity of $H$ that
@@
while reorganizing the term $\mathcal{B}^{n}$ gives
@@
$\phantom{xxxxx}\;+\lambda^{2}\sum_{\sigma\in\mathcal{E}^{D}}a_{\sigma}\phi_{\sigma}^{D}(\phi_{K}^{n}-\phi_{K}^{n-1}).$
Then using the elementary convexity inequality $a(a-b)\geq(a^{2}-b^{2})/2$ in the above term and combining the result with (45) in (44) provides the desired result (42). ∎
@@ 3.3. Further uniform estimates on the discrete solution
To pass to the limit in the scheme and to identify the limit as a weak solution, we need to extract some further uniform estimates, in particular the discrete $L^{2}_{\text{loc}}(H^{1})$ estimates on the discrete counterparts of $u_{0}$ and $\sqrt{u_{i}u_{0}}$. We prove these estimates in Lemma 3.7. As an intermediate result we need a uniform bound on the discrete free energy.
Lemma 3.6.
@@
Because of the bound $0\leq u_{i,K}^{n}\leq 1$ for all $i$ and $K$, it is clear that the first two contributions of (41) remain uniformly bounded by a constant depending only on $\Omega$. Concerning the last contribution observe that if one defines $\phi_{K}^{D}$ and $\phi_{\sigma}^{D}$ as the averages of $\phi^{D}$ on $K\in\mathcal{T}$ and $\sigma\in\mathcal{E}$ respectively, then a suitable reorganization of the sum shows that
@@
Then since $f$ and $U$ are bounded, we deduce from (46) that there exists $C$ depending only on $\|f\|_{\infty}$, $\max_{i}|z_{i}|$, $\Omega$, $\|\phi^{D}\|_{H^{1}}$, $\lambda$ and $\zeta_{\mathcal{T}}$ such that
@@
Next lemma shows that the control of the energy dissipation $\sum_{n\geq 1}\tau^{n}\mathcal{D}_{\mathcal{T}}^{n}$ inferred from Proposition 3.5 gives some $L^{2}_{\text{loc}}(H^{1})$ type control on the discrete counterparts of $u_{0}$, $\sqrt{u_{0}}$ and $\sqrt{u_{i}u_{0}}$.
Lemma 3.7.
There exists $C>0$ depending only on $\Omega$, $\phi^{D}$, $\lambda$, $f$, ${(z_{i})}_{i}$, ${(D_{i})}_{i}$ and $\zeta_{\mathcal{T}}$ such that, for all $N\geq 1$, there holds
(47) $\sum_{n=1}^{N}\tau^{n}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}\\ +\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}\\ +\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(u_{0,K}^{n}-u_{0,L}^{n}\right)^{2}\leq C(1+t^{N}).$
@@
(48) $\sum_{n=1}^{N}\tau^{n}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}-u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)^{2}\leq C(1+t^{N})$
@@
(49) $\sum_{n=1}^{N}\tau^{n}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(u_{i,K}^{n}u_{0,K}^{n}-u_{i,L}^{n}u_{0,L}^{n}\right)^{2}\leq C(1+t^{N}).$
@@
for some $\kappa>0$ uniform w.r.t. $K$, $i$ and $n$. As a consequence, using furthermore that $(a+b)^{2}\geq\frac{1}{2}a^{2}-b^{2}$,
$\mathcal{D}_{i,\sigma}^{n}\geq\kappa a_{\sigma}\cosh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right)\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}-\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\\ -\kappa a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}+\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\sinh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right).$
Since $|\phi_{K}^{n}|\leq C$ owing to (39), one has $\sinh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right)\leq C(\phi_{K}^{n}-\phi_{L}^{n})^{2}$. Using moreover that $10<u_{i,K}^{n},u_{0,K}^{n}<$ 1 and that $\cosh(a)\geq 1$, one gets that
@@
$\mathcal{D}_{\mathcal{T}}^{n}\geq\kappa\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}\\ +\kappa\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(u_{0,K}^{n}-u_{0,L}^{n}\right)^{2}-C\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}(\phi_{K}^{n}-\phi_{K\sigma}^{n})^{2}.$
@@
(51) $\mathcal{D}_{\mathcal{T}}^{n}\geq\kappa\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}+\kappa\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(u_{0,K}^{n}-u_{0,L}^{n}\right)^{2}-C.$
Moreover, the inequality $\sum^{I}_{i=0}\sqrt{u_{i,K}^{n}u_{i,L}^{n}}\leq 1$ gives that
$\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}\\ \geq\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left((1-u_{0,K}^{n})u_{0,K}^{n}+(1-u_{0,L}^{n})u_{0,L}^{n}-2(1-\sqrt{u_{0,K}^{n}u_{0,L}^{n}})\sqrt{u_{0,K}^{n}u_{0,L}^{n}}\right)\\ =\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}-\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left({u_{0,K}^{n}}-{u_{0,L}^{n}}\right)^{2},$
@@
$\mathcal{D}_{\mathcal{T}}^{n}\geq\kappa\sum_{\sigma\in\mathcal{E}_{\text{int}}}a_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}-C.$
To deduce (47), it eventually remains to remark from (42) and Lemma 3.6 that there exists $C$ depending neither on $h$, $\Delta t$, $N$ nor on the initial data $U^{0}=\left(u_{i}^{0}\right)_{1\leq i\leq I}$ (provided it fulfills (7)) such that $\sum_{n=1}^{N}\tau^{n}\mathcal{D}_{\mathcal{T}}^{n}\leq C.$ Combining this with (51) yields (47).
@@
[…]
which again, after summation on $n=1,\dots,N$ and $\sigma\in\mathcal{E}_{\text{int}}$, can be controlled thanks to (47).
@@
Summing over $n=1,\dots,N$ and $\sigma\in\mathcal{E}_{\text{int}}$ and using (47) gives the desired result. ∎
One also deduces the following discrete $L^{2}_{\text{loc}}(L^{2})^{d}$ estimates on the fluxes, which amount to some discrete $L^{2}_{\text{loc}}(H^{1})^{\prime}$ estimate on time increments of the discrete counterpart to $\partial_{t}u_{i}$.
Lemma 3.8.
There exists $C$ depending only on $\Omega$, $\phi^{D}$, $\lambda$, $f$, ${(z_{i})}_{i}$, ${(D_{i})}_{i}$ and $\zeta_{\mathcal{T}}$ such that
(52) $\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}\frac{d_{\sigma}}{m_{\sigma}}\left|F_{i,K\sigma}^{n}\right|^{2}\leq C(1+t^{N}).$
@@
To obtain discrete $L^{2}_{\text{loc}}(L^{2})^{d}$ estimates on the fluxes, we need to exploit the discrete uniform estimates we have for $(U_{K}^{n},\phi_{K}^{n})_{K\in\mathcal{T},n\geq 1}.$ In this regard, we rewrite the fluxes in a different way before taking advantage of some properties of the Bernoulli function (28). One splits the flux (2) into two parts corresponding to convection and diffusion respectively:
@@
The flux $(F_{i,K\sigma}^{n})_{\sigma,n}$ is bounded in $L^{2}_{\text{loc}}(L^{2})^{d}$ in the sense of (52) if both $(F_{i,K\sigma}^{\text{conv},n})_{\sigma,n}$ and $(F_{i,K\sigma}^{\text{diff},n})_{\sigma,n}$ are.
@@
The $L^{2}_{\text{loc}}(L^{2})^{d}$ character of the above expression directly follows from the uniform bound on $u_{i,K}^{n}$, $0\leq i\leq I$ and from the discrete $L^{\infty}(H^{1})$ bound on $(\phi_{K}^{n})_{K,n}$ stated in Proposition 3.3. Therefore,
$\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\text{int}}}\frac{d_{\sigma}}{m_{\sigma}}\left|F_{i,K\sigma}^{\text{conv},n}\right|^{2}\leq Ct^{N}.$
@@
From the discrete $L^{\infty}(H^{1})$ bound on $(\phi_{K}^{n})_{K,n}$, one can uniformly estimate the reminder. For the other term, since $u_{i,K}^{n}u_{0,L}^{n}-u_{i,L}^{n}u_{0,K}^{n}=u_{i,K}^{n}u_{0,K}^{n}-u_{i,L}^{n}u_{0,L}^{n}+(u_{i,K}^{n}+u_{i,L}^{n})(u_{0,L}^{n}-u_{0,K}^{n}),$ then

### v2

@@ 3.2. Energy dissipation at the discrete level
The next proposition is about the thermodynamical consistency of our scheme and the decay of a discrete counterpart of the free energy.
Proposition 3.6 (Free energy decay).
@@
(41) $\mathcal{H}_{\mathcal{T}}^{n}=\sum_{K\in\mathcal{T}}m_{K}H(U_{K}^{n})+\frac{\lambda^{2}}{2}\sum_{\sigma\in\mathcal{E}}a_{\sigma}(\phi_{K}^{n}-\phi_{K\sigma}^{n})^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\lambda^{2}\sum_{\sigma\in\mathcal{E}^{D}}a_{\sigma}\phi_{\sigma}^{D}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\phi_{\sigma}^{D}-\phi_{K}^{n})},$
@@
$\mathcal{D}_{\mathcal{T}}^{n}=\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}F_{i,K\sigma}^{n}(\mu_{i,K}^{n}-\mu_{i,L}^{n})$
@@
(43) $\mathcal{D}_{i,\sigma}^{n}:=F_{i,K\sigma}^{n}(\mu_{i,K}^{n}-\mu_{i,L}^{n})\geq 0,\qquad\forall i=1,\dots,I,\;\sigma=K|L\in\mathcal{E}_{{\rm int}},$
whence the non-negativity of $\mathcal{D}^{n}$. As $u_{0,K}^{n}>0$ for all $K\in\mathcal{T}$ and $n\geq 1$, and as $y\mapsto e^{y}$ is strictly increasing, one gets that $\mathcal{D}_{i,\sigma}^{n}=0$ iff $\mu_{i,K}^{n}=\mu_{i,L}^{n}$.
To proceed, we first define by $\mu_{i,K}^{n}=\log\left(\frac{u_{i,K}^{n}}{u_{0,K}^{n}}\right)+z_{i}\phi_{K}^{n}=\log(w_{i,K}^{n})$ the electrochemical potential of species $i$, then we multiply the discrete conservation law (25) by $\tau^{n}\mu_{i,K}^{n}$ and sum over $i=1,\dots,I$ and $K\in\mathcal{T}$. Thanks to the local conservativity of the fluxes, i.e., $F_{i,K\sigma}^{n}+F_{i,L\sigma}^{n}=0$ for all $\sigma=K|L\in\mathcal{E}_{\rm int}$, and to the no-flux boundary condition $F_{i,K\sigma}^{n}=0$ for $\sigma\in\mathcal{E}_{\text{ext}}$, $i=1,\dots,I$, a discrete integration by parts in space allows the flux terms to be rewritten as a sum over interior edges only, namely $\sum_{K\in\mathcal{T}}\sum_{\sigma\in\mathcal{E}_{K}}F_{i,K\sigma}^{n}\mu_{i,K}^{n}=\sum_{\sigma\in\mathcal{E}_{\rm int}}\mathcal{D}_{i,\sigma}^{n}$ for all $i=1,\dots,I$. This leads to the identity
(44) $\mathcal{A}_{\mathcal{T}}^{n}+\mathcal{B}_{\mathcal{T}}^{n}+\tau^{n}\mathcal{D}_{\mathcal{T}}^{n}=0.$
@@
[…]
where $\widetilde{H}:\left\{(u_{i})_{0\leq i\leq I}\;\middle|\;\sum_{i=0}^{I}u_{i}=1\right\}\to[-\log(I+1),+\infty)$, defined by $\widetilde{H}(u_{0},\dots,u_{I}):=\sum_{i=0}^{I}\log(u_{i})$, and
@@
[…]
From the convexity of $\widetilde{H}$ and the identity $\widetilde{H}(u_{0},\dots,u_{I})=H(U)$, we deduce that
@@
while the term $\mathcal{B}^{n}_{\mathcal{T}}$ can be rewritten by observing that each interior edge is counted twice in the sum $\sum_{K\in\mathcal{T}}\sum_{\sigma\in\mathcal{E}_{K}}$. The resulting equivalent reformulation is
@@
$\phantom{xxxxx}\;+\lambda^{2}\sum_{\sigma\in\mathcal{E}^{D}}a_{\sigma}\phi_{\sigma}^{D}(\phi_{K}^{n}-\phi_{K}^{n-1}),$
which is valid since boundary edges $\sigma\in\mathcal{E}^{N}$ do not contribute to the first sum, and the contributions from boundary edges $\sigma\in\mathcal{E}^{D}$ cancel with the second part of the new expression of $\mathcal{B}_{\mathcal{T}}^{n}$. Then using the elementary convexity inequality $a(a-b)\geq(a^{2}-b^{2})/2$ in the above term and combining the result with (45) in (44) provides the desired result (42). ∎
@@ 3.3. Further uniform estimates on the discrete solution
To pass to the limit in the scheme and to identify the limit as a weak solution, we need to extract some further uniform estimates, in particular the discrete $L^{2}_{\rm loc}(H^{1})$ estimates on the discrete counterparts of $u_{0}$ and $\sqrt{u_{i}u_{0}}$. We prove these estimates in Lemma 3.8. As an intermediate result we need a uniform bound on the discrete free energy.
Lemma 3.7 (Uniform bound for the free energy).
@@
Because of the bound $0\leq u_{i,K}^{n}\leq 1$ for all $i$ and $K$, it is clear that the first two contributions of (41) remain uniformly bounded by a constant depending only on $\Omega$. Indeed, for the first term, this is a direct consequence of the definition (10) of $H$, whereas the second term can be controlled thanks to Proposition 3.3. Concerning the last contribution observe that if one defines $\phi_{K}^{D}$ and $\phi_{\sigma}^{D}$ as the averages of $\phi^{D}$ on $K\in\mathcal{T}$ and $\sigma\in\mathcal{E}$ respectively, then a suitable reorganization of the sum shows that
@@
Then since $f$ and $U$ are bounded, we deduce from (46) that there exists $C$ depending only on $\|f\|_{\infty}$, $\max_{i}|z_{i}|$, $\Omega$, $\|\phi^{D}\|_{H^{1}}$, $\lambda$, and $\zeta_{\mathcal{T}}$ such that
@@
The next lemma shows that the control of the energy dissipation $\sum_{n\geq 1}\tau^{n}\mathcal{D}_{\mathcal{T}}^{n}$ inferred from Proposition 3.6 gives some $L^{2}_{\rm loc}(H^{1})$ type control on the discrete counterparts of $u_{0}$, $\sqrt{u_{0}}$, and $\sqrt{u_{i}u_{0}}$.
Lemma 3.8 (Uniform bounds on $u_{0}$, $\sqrt{u_{0}}$, and $\sqrt{u_{i}u_{0}}$).
There exists $C>0$ depending only on $\Omega$, $\phi^{D}$, $\lambda$, $f$, ${(z_{i})}_{i}$, ${(D_{i})}_{i}$, and $\zeta_{\mathcal{T}}$ such that, for all $N\geq 1$, there holds
(47) $\sum_{n=1}^{N}\tau^{n}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}\\ +\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}\\ +\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{0,K}^{n}-u_{0,L}^{n}\right)^{2}\leq C(1+t^{N}).$
@@
(48) $\sum_{n=1}^{N}\tau^{n}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{i,K}^{n}\sqrt{u_{0,K}^{n}}-u_{i,L}^{n}\sqrt{u_{0,L}^{n}}\right)^{2}\leq C(1+t^{N})$
@@
(49) $\sum_{n=1}^{N}\tau^{n}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{i,K}^{n}u_{0,K}^{n}-u_{i,L}^{n}u_{0,L}^{n}\right)^{2}\leq C(1+t^{N}).$
@@
for some $\kappa>0$ uniform w.r.t. $K$, $i$, and $n$. As a consequence, using furthermore that $(a+b)^{2}\geq\frac{1}{2}a^{2}-b^{2}$,
$\mathcal{D}_{i,\sigma}^{n}\geq\kappa a_{\sigma}\cosh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right)\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}-\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\\ -{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}\kappa a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,L}^{n}}+\sqrt{u_{i,L}^{n}u_{0,K}^{n}}\right)^{2}\sinh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right).$
Since $|\phi_{K}^{n}|\leq C$ owing to (39), one has $\sinh^{2}\left(\frac{z_{i}}{4}(\phi_{K}^{n}-\phi_{L}^{n})\right)\leq C(\phi_{K}^{n}-\phi_{L}^{n})^{2}$. Using moreover that $0<u_{i,K}^{n},u_{0,K}^{n}<1$ and that $\cosh(a)\geq 1$, one gets that
@@
$\mathcal{D}_{\mathcal{T}}^{n}\geq\kappa\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}\\ +\kappa\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{0,K}^{n}-u_{0,L}^{n}\right)^{2}-C\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}(\phi_{K}^{n}-\phi_{K\sigma}^{n})^{2}.$
@@
(51) $\mathcal{D}_{\mathcal{T}}^{n}\geq\kappa\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}+\kappa\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(u_{0,K}^{n}-u_{0,L}^{n}\right)^{2}-C.$
Moreover, since $\sqrt{ab}\leq(a+b)/2$ for all $a,b\geq 0$, there holds
$\sum^{I}_{i=0}\sqrt{u_{i,K}^{n}u_{i,L}^{n}}\leq\sum^{I}_{i=0}\frac{u_{i,K}^{n}+u_{i,L}^{n}}{2}=1,\qquad\forall\sigma=K|L\in\mathcal{E}_{\text{int}}.$
This provides
$\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{i,K}^{n}u_{0,K}^{n}}-\sqrt{u_{i,L}^{n}u_{0,L}^{n}}\right)^{2}\\ \geq\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left((1-u_{0,K}^{n})u_{0,K}^{n}+(1-u_{0,L}^{n})u_{0,L}^{n}-2(1-\sqrt{u_{0,K}^{n}u_{0,L}^{n}})\sqrt{u_{0,K}^{n}u_{0,L}^{n}}\right)\\ =\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}-\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left({u_{0,K}^{n}}-{u_{0,L}^{n}}\right)^{2},$
@@
$\mathcal{D}_{\mathcal{T}}^{n}\geq\kappa\sum_{\sigma\in\mathcal{E}_{\rm int}}a_{\sigma}\left(\sqrt{u_{0,K}^{n}}-\sqrt{u_{0,L}^{n}}\right)^{2}-C.$
To deduce (47), it eventually remains to remark from (42) and Lemma 3.7 that there exists $C$ depending on none of $h$, $\Delta t$, $N$ nor on the initial data $U^{0}=\left(u_{i}^{0}\right)_{1\leq i\leq I}$ (provided it fulfills (7)) such that $\sum_{n=1}^{N}\tau^{n}\mathcal{D}_{\mathcal{T}}^{n}\leq C.$ Combining this with (51) yields (47).
@@
$\leq\;$ $\sum_{i=1}^{I}\Big{(}\sqrt{u^{n}_{0,K}u^{n}_{i,K}}-\sqrt{u^{n}_{0,L}u^{n}_{i,L}}\Big{)}^{2}+(u_{0,K}^{n}-u_{0,L}^{n})^{2},$
which again, after summation on $n=1,\dots,N$ and $\sigma\in\mathcal{E}_{\rm int}$, can be controlled thanks to (47).
@@
Summing over $n=1,\dots,N$ and $\sigma\in\mathcal{E}_{\rm int}$ and using (47) gives the desired result. ∎
One also deduces the following discrete $L^{2}_{\rm loc}(L^{2})^{d}$ estimates on the fluxes, which amount to some discrete $L^{2}_{\rm loc}(H^{1})^{\prime}$ estimate on time increments of the discrete counterpart to $\partial_{t}u_{i}$.
Lemma 3.9 (Uniform bounds on the fluxes).
There exists $C$ depending only on $\Omega$, $\phi^{D}$, $\lambda$, $f$, ${(z_{i})}_{i}$, ${(D_{i})}_{i}$, and $\zeta_{\mathcal{T}}$ such that
(52) $\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}\frac{d_{\sigma}}{m_{\sigma}}\left|F_{i,K\sigma}^{n}\right|^{2}\leq C(1+t^{N}).$
@@
To obtain discrete $L^{2}_{\rm loc}(L^{2})^{d}$ estimates on the fluxes, we need to exploit the discrete uniform estimates we have for $(U_{K}^{n},\phi_{K}^{n})_{K\in\mathcal{T},n\geq 1}.$ In this regard, we rewrite the fluxes in a different way before taking advantage of some properties of the Bernoulli function (28). One splits the flux (26) into two parts corresponding to convection and diffusion respectively:
@@
The flux $(F_{i,K\sigma}^{n})_{\sigma,n}$ is bounded in $L^{2}_{\rm loc}(L^{2})^{d}$ in the sense of (52) if both $(F_{i,K\sigma}^{\text{conv},n})_{\sigma,n}$ and $(F_{i,K\sigma}^{\text{diff},n})_{\sigma,n}$ are.
@@
The $L^{2}_{\rm loc}(L^{2})^{d}$ character of the above expression directly follows from the uniform bound on $u_{i,K}^{n}$, $0\leq i\leq I$ and from the discrete $L^{\infty}(H^{1})$ bound on $(\phi_{K}^{n})_{K,n}$ stated in Proposition 3.3. Therefore,
$\sum_{i=1}^{I}\sum_{n=1}^{N}\tau^{n}\sum_{\sigma\in\mathcal{E}_{\rm int}}\frac{d_{\sigma}}{m_{\sigma}}\left|F_{i,K\sigma}^{\text{conv},n}\right|^{2}\leq Ct^{N}.$
@@
From the discrete $L^{\infty}(H^{1})$ bound on $(\phi_{K}^{n})_{K,n}$, one can uniformly estimate the remainder. For the other term, since $u_{i,K}^{n}u_{0,L}^{n}-u_{i,L}^{n}u_{0,K}^{n}=u_{i,K}^{n}u_{0,K}^{n}-u_{i,L}^{n}u_{0,L}^{n}+(u_{i,K}^{n}+u_{i,L}^{n})(u_{0,L}^{n}-u_{0,K}^{n}),$ then